Grade 12th passPhysical Chemistry

if 240 g of carbon is taken in a container to convert completely to co2 but in industry it has been found that 280 g of co was also formed along with co2. find the mole percentage yield of co2.​

bittu kumar
6 Years agoGrade 12th pass

3 Answers

Arun
6 Years ago
Yield of CO2 = (Amount of CO2 actually formed) / (Amount of CO2 that can be formed by complete oxidation of given carbon)
 
C + O2 \rightarrow CO2
 
2C + O2 \rightarrow 2CO
 
1 mole of Carbon(C) yields 1 mole of CO2.
 
Number of moles of carbon = (240)/12 = 20 moles.
 
Number of moles of carbon that is used to form CO = Number of moles of CO formed = (280)/(28) = 10 moles
 
\implies Number of moles of carbon used in the formation of CO2 = Number of moles of CO2 formed = 20 – 10 = 10 moles
 
Number of moles of CO2 that can be formed by the complete oxidation of given amount of carbon = Number of moles of carbon = 20 moles
 
\implies Yield of CO2 = (10)/(20) = 0.5
 
\impliespercentage yield of CO2 = 0.5 * 100% = 50%
 
 
Vikas TU
6 Years ago
C + O2 \rightarrow CO2
2C + O2 \rightarrow 2CO
1 mole of Carbon(C) yields 1 mole of CO2.
Number of moles of carbon = (240)/12 = 20 moles.
Number of moles of carbon that is used to form CO = Number of moles of CO formed = (280)/(28) = 10 moles
\implies Number of moles of carbon used in the formation of CO2 = Number of moles of CO2 formed = 20 – 10 = 10 moles
Number of moles of CO2 that can be formed by the complete oxidation of given amount of carbon = Number of moles of carbon = 20 moles
\implies Yield of CO2 = (10)/(20) = 0.5
\impliespercentage yield of CO2 = 0.5 * 100% = 50%
 
MANTHAN RAO
6 Years ago
ANSWER 50%
Explanation:
Yield of CO2 = (Amount of CO2 actually formed) / (Amount of CO2 that can be formed by complete oxidation of given carbon)
C + O2 \rightarrow CO2
2C + O2 \rightarrow 2CO
1 mole of Carbon(C) yields 1 mole of CO2.
Number of moles of carbon = (240)/12 = 20 moles.
Number of moles of carbon that is used to form CO = Number of moles of CO formed = (280)/(28) = 10 moles
\implies Number of moles of carbon used in the formation of CO2 = Number of moles of CO2 formed = 20 – 10 = 10 moles
Number of moles of CO2 that can be formed by the complete oxidation of given amount of carbon = Number of moles of carbon = 20 moles
\implies Yield of CO2 = (10)/(20) = 0.5
\impliespercentage yield of CO2 = 0.5 * 100% = 50%